# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 9. Conclusion
LLMs lend themselves as strong tools for comprehending complex linguistic patterns and generating logical and contextually coherent responses. However, such powerful models also entail potential privacy and security risks. In this survey, we first provided a detailed overview of LLMs’ security and privacy challenges. We then discussed existing mitigation and defense strategies against these security attacks and privacy attacks, as well as highlighting their strengths and limitations.
In our investigation, we found that LLMs are highly vulnerable to the attacks discussed in the corresponding sections. According to our survey, there are a limited number of mitigation techniques to prevent those attacks against LLMs. The existing mitigation techniques that are applicable to relatively small LMs could potentially be used for LLMs. However, extensive research studies should be performed to evaluate and tailor the existing solutions to LLMs. Based on our analysis, we also outlined future research directions focusing on security and privacy aspects pointing out key research gaps, and illustrating open research problems. The overarching goal is to enhance the reliability and utility of LLMs through comprehensive exploration and resolution of these vulnerabilities and offer pathways for future research toward secure and privacy-preserving LLM systems.
Acknowledgements.
This work is partially supported by the U.S. Department of Homeland Security Grant Award Number 17STCIN00001-05-00. Further, M. Hadi Amini’s work is partly supported by the U.S. Department of Homeland Security under Grant Award Number 23STSLA00016-01-00. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security.

### v2

@@ 9. Conclusion
LLMs lend themselves as strong tools for comprehending complex linguistic patterns and generating logical and contextually coherent responses. However, such powerful models also entail potential privacy and security risks. In this survey, we first provided a detailed overview of LLMs’ security and privacy challenges. We then discussed and analyzed the LLM vulnerabilities from both security and privacy aspects, existing mitigation and defense strategies against these security attacks and privacy attacks, as well as highlighted their strengths and limitations. In our investigation, we found that LLMs are highly vulnerable to the discussed attacks. According to our survey, there are a limited number of mitigation techniques to prevent those attacks against LLMs. The existing mitigation techniques that are applicable to relatively small LMs could potentially be used for LLMs. However, extensive research studies should be performed to evaluate and tailor the existing solutions to LLMs. Based on our analysis, we also outlined future research directions focusing on security and privacy aspects, pointed out key research gaps, and illustrated open research problems. The overarching goal is to enhance the reliability and utility of LLMs through comprehensive exploration and resolution of these vulnerabilities and offer pathways for future research toward secure and privacy-preserving LLM systems.
ACKNOWLEDGEMENTS: This work is partially supported by the U.S. Department of Homeland Security Grant Award Number 17STCIN00001-05-00. Further, M. Hadi Amini’s work is partly supported by the U.S. Department of Homeland Security under Grant Award Number 23STSLA00016-01-00. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security.